if work done equation is given as W = 1/2 MV2 - 1/2 MU2 show that equation is dimensionally correct

1
Expert's answer
2021-04-05T11:09:14-0400

Let's show that this equation is dimensionally correct:


"W=\dfrac{1}{2}MV^2-\dfrac{1}{2}MU^2,""ML^{2}T^{-2}=M(LT^{-1})^2-M(LT^{-1})^2,""ML^{2}T^{-2}=ML^{2}T^{-2}."

Therefore, this equation is dimensionally correct.